CAN SOMEONE PLEASE HELP ME WITH THIS PROBLEM:
The length a spring is stretched from its natural length with work, W foot-pounds, given by L = checksign (2W/k) where k is a constant for the given spring. If a certain spring has a constant of 76.4, and the spring is to be stretched 3.6 feet from its natural length, how much work will be necessary?

Answer by scott8148(6628)

squaring __ L^2=2W/k __ multiplying by k and dividing by 2 __ (kL^2)/2=W

76.4 * 3.6^2 /2 =W __ 495=W (approx)